OpenStudy (anonymous):

Solve the system algebraically. 2x + y - 10 = 0 x - y - 5 = 0 What is the value of y? -5/3 0 5

OpenStudy (jdoe0001):

just add them up, vertically

OpenStudy (ivettef365):

you can isolate x and you can said that x=y+5 now on the first equation substitute x for (y+5) and find value of y

OpenStudy (jdoe0001):

"y" drops right off, you'd get "x", then use your "x" value on either equation

OpenStudy (ivettef365):

yes that is another way, you can add

OpenStudy (anonymous):

so what would the answer be i did that and got like thirteen>

OpenStudy (ivettef365):

you have 2x +y -10 = 0 x -y -5 = 0 ---------------- 3x -15 = 0

OpenStudy (ivettef365):

solve for x

OpenStudy (anonymous):

oooooooooooooooohhhhhhhhhhhhhhhhhhhhhhhhhhhh. thank you so much. i am doing an algebra class to get ahead this summer so it is like stepping into a whole new world. i thank you for that help @ivettef365

OpenStudy (ivettef365):

so what did you get for x

OpenStudy (anonymous):

...3x = 15

OpenStudy (anonymous):

you also couldve doubled the bottom one in order to cancel the x variable. it wouldve become 2x - 2y - 10

OpenStudy (anonymous):

subtracting that from the top equation would give 3y = 0